OpenStudy (anonymous):

HOw many solutions does this equation have? 9/x-5+4/x^2-3x-10=6/x+2

jhonyy9 (jhonyy9):

x2 -3x -10=(x-5)(x+2)

jhonyy9 (jhonyy9):

9(x+2)+4=6(x-5) 9x+18+4=6x-30 3x=-52 x=52/3